# Camouflaged Object Tracking: A Benchmark

### v1

@@ Abstract
1 Introduction
@@
2 Related Works
2.1 Visual Tracking Algorithms
DCF-based trackers treat visual tracking as an online regression problem. The MOSSE filter [3] was an early and notable example. Afterwards, enhancements such as scale filters and regularization techniques to improve robustness have been introduced [17, 45, 16, 42, 44]. However, these trackers still struggle with complex conditions due to their reliance on handcrafted features. To overcome these challenges, deep learning-based trackers have become increasingly popular. These trackers use neural networks to automatically learn features. A key innovation in this area was SiamFC [2], which utilized Siamese networks [14] for visual tracking, learning a similarity function to generate response maps for the search area. Subsequent methods, such as SiamRPN++ [41], SiamR-CNN [62], SiamCAR [28], and Ocean [85], have further enhanced accuracy and robustness. Recently, Vision Transformers (ViTs) [19] have introduced new possibilities for visual tracking. ViTs offer potential for streamlining and unifying tracking frameworks [19, 51, 69, 78]. For instance, OStrack [78] combines feature extraction with relational modeling for efficient single-stream tracking, HIPTrack [6] introduces a history prompting network that leverages past object masks and visual features for more precise guidance. However, most current research focuses on tracking generic objects, with less attention given to more specialized scenarios, particularly the tracking of camouflaged objects. In this work, we introduce a dataset for tracking camouflaged objects and enhance HIPTrack with multi-level features to address this challenge.
2.2 Visual Tracking Benchmarks
@@
Currently, there are two main types of benchmark datasets: general benchmarks and specific benchmarks [24]. 1) General Benchmarks: OTB [70, 72] and VOT [61, 39, 36, 35, 37, 38] are among the initial platforms for unified testing data and evaluation standards for general visual tracking. NFS [33] focuses on high-frame-rate videos. Many large-scale benchmarks have been introduced for training deep learning-based trackers. For example, TrackingNet [58] aims to provide resources for training and evaluating general object tracking algorithms, containing over 30,000 sequences. GOT-10K [32] includes over 10,000 videos from the WordNet semantic hierarchy to develop the generalization capability of trackers. LaSOT [23] encompasses 1,400 long-term video sequences, with an additional 150 sequences and a new evaluation scheme. 2) Specific Benchmarks: UAV123 [1] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [52] and PTB [64] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [24] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [86] has been introduced for tracking small and fast-moving objects, while TRO focuses on tracking reflected objects.
@@
2.3 Dealing with camouflaged Objects in Vision
Camouflaged objects are difficult to distinguish from their backgrounds because they blend into their surroundings. Unlike regular objects, they have similar textures, colors, and patterns as their environment, creating visual ambiguity. Recently, many computer vision techniques have been developed to tackle these challenges. For instance, Fan et al. [20] introduced the task of Camouflaged Object Detection (COD) and established a dataset COD10K with 10,000 images, along with a simple yet effective framework SINet for efficient and precise detection of camouflaged objects. Liu et al. [48] proposed a scale-adaptive correlation filter algorithm that integrates FHOG and LBP features for Camouflaged Object Segmentation (COS). Cheng et al. [13] developed a Video Camouflaged Object Detection (VCOD) framework that captures inter-frame motion and uses a spatio-temporal transformer for consistency, along with the MoCA-Mask dataset providing detailed pixel-level annotations. Hu et al. [31] tackled visual camouflage challenges in UAV surveillance with a novel trajectory planning control framework. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on consistently identifying camouflaged objects in videos while adapting to changes in position, scale, and appearance, remains under-explored. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), specifically designed to support the development and evaluation of algorithms tailored for tracking camouflaged objects.
3 Construction of the COTD Dataset
3.1 Data Collection
@@
In the video selection process, we apply stringent criteria to ensure the high quality and practicality of the data. Specifically, videos are evaluated and selected based on three main criteria: 1) the diversity and complexity of the environment, 2) the frequency with which targets cross frame boundaries, and 3) the overall length and continuity of the videos. We prioritize videos that present additional challenges, such as a high degree of similarity between targets and backgrounds. Due to resource constraints in collecting, processing, and annotating video data, we select 200 sequences for the benchmark test. This ensures that each video is of high quality with comprehensive annotations, adhering to our available time, budget, and manpower constraints. The COTD dataset comprises 20 categories, namely: bird, bug, butterfly, cat, crab, dog, fish, frog, grasshopper, leopard, lizard, mantis, octopus, person, phasmatodea, seahorse, rabbit, shrimp, snake, and spider. To visually represent the composition of the dataset, we create a chart illustrating the number of sequences per category. Additionally, we generate a series of graphs depicting the maximum, minimum, and average number of frames for each category. These visual aids not only enhance the understanding of the dataset’s diversity but also highlight the challenges in tracking camouflaged objects across different scenarios. Fig. 2 presents the distribution of sequences across the 20 categories, while Fig. 3 provides a detailed view of the frame characteristics for each category, such as the range and average number of frames. These figures are crucial in demonstrating the dataset’s comprehensiveness and its potential as a robust benchmark for evaluating the performance of camouflaged object tracking algorithms. Additionally, to highlight the challenges associated with tracking camouflaged objects, we evaluate seven state-of-the-art trackers on the COTD dataset and compare them with two popular generic object tracking benchmarks. The AUC results are presented in Table 1. Notably, the AUC of these trackers on the TrackingNet benchmark is more than 15.0% higher than on the COTD dataset, underscoring the significant challenges posed by tracking camouflaged objects.
@@
Table 1: AUC (%) Comparison of state-of-the-art trackers on the COTD and generic object tracking benchmarks.
@@
3.2 Data Annotation
We follow the principles in [58] for sequence annotation. Each video frame is meticulously annotated by a team of expert annotators, specifically students engage in visual tracking research. Annotators draw and edit axis-aligned bounding boxes on each frame to represent the most compact boundary of the object. Visible targets are annotated as present, while those missing, out of view, or occluded are marked accordingly. To enhance the efficiency and quality of the annotations, we adopt a stringent manual annotation process. Each frame’s annotation underwent thorough review to ensure the accuracy and consistency of the bounding boxes. Annotators meticulously label each object frame-by-frame, ensuring accurate identification even within complex backgrounds. To further ensure high-quality annotations, we implement a double-check mechanism where a second annotator review the initial annotations to resolve any discrepancies. All differences are addressed through discussion and consensus. This rigorous manual annotation process and double-check mechanism guarantee the high quality of the dataset. Examples of bounding box annotations in the COTD are shown in Fig. 4.
3.3 Attributes
Figure 5: Distribution of sequences with each attribute in COTD.
Following the prevalent tracking benchmark methodologies [93], [32], and [72], we meticulously annotate each video sequence with detailed attributes to enable an in-depth analysis of tracking algorithms’ performance. The proposed COTD dataset features 12 attributes, with each sequence including Background Clutter (BC) to highlight the dataset’s camouflage characteristics. In addition, we annotate the following 11 attributes: (1) Illumination Variation (IV), (2) Scale Variation (SV), (3) Deformation (DEF), (4) Motion Blur (MB), (5) Fast Motion (FM), which is marked when the object center moves by at least 20% of its own size between consecutive frames, (6) Out-of-View (OV), (7) Low Resolution (LR), marked when the object area is smaller than 900 pixels, (8) Partial Occlusion (POC), (9) Rotation (ROT), (10) Full Occlusion (FOC), and (11) Aspect Ratio Change (ARC), marked when the object’s bounding box aspect ratio falls outside the range [0.5, 2]. Fig.5 shows the distribution of these attributes on COTD. As can be seen, the most common challenge in COTD is Motion Blur. In addition, the Deformation and Partial Occlusion also present frequently in COTD. By comprehensively considering these attributes, we can conduct a thorough evaluation of tracking algorithms under various challenging conditions.
4 Method
4.1 Overall Architecture
Figure 7: Structure of the Multi-Level Feature Sensing Encoder.
4.2 Multi-Level Feature Sensing Encoder
5 Evaluation
We use one-pass evaluation (OPE) to assess precision and success rate, following the approach outlined in [22, 57]. Precision is the pixel distance between the centers of the estimated and ground truth bounding boxes, while success rate is based on the intersection over union (IoU) exceeding 0.5 [71, 22, 57]. Trackers are ranked using precision at 20 pixels (PRC) and the area under the curve (AUC) of the success plot.
5.1 Trackers for Comparison
5.2 Evaluation Results
Attribute-based performance: We conducted a comprehensive performance evaluation on eleven common attributes to better analyze and understand the performance of different trackers on the COTD dataset. HIPTrack-MLS consistently ranks first across most attribute subsets. Due to paper length constraints, we present the PRC results for four common challenges in Fig. 9, including Aspect Ration Change (ARC), Illumination Variation (IV), Out-of-View (OV), and Partial Occlusion (POC). More results can be found in the supplementary material. As shown, HIPTrack-MLS consistently ranks first across all these attribute subsets, surpassing the baseline HIPTrack by at least 1.0% in all these subsets. Notably, HIPTrack-MLS achieves a precision of 0.271 in the IV subset, surpassing the second place HIPTrack by 3.4%. For the ARC and OV attributes, HIPTrack-MLS shows improvements of 1.3% and 1.0%, respectively, over HIPTrack. These results highlight the effectiveness of the proposed method for tracking camouflaged objects.
Figure 10: Qualitative evaluation on 3 sequences from COTD, i.e. mantis-2, person-9, and phasmatodea-10 from top to bottom. The results of different methods have been shown with different colors, and ‘GT’ denotes the ground truth.
Qualitative Evaluation: In Fig. 10, we demonstrate qualitative comparisons of our method with 9 top trackers, including HIPTrack [5], DropTrack [69], DCPT [92], AQATrack [75], ARTrack [68], GRM [27], ROMTrack [7], SeqTrack [11], and ZoomTrack [34]. These examples face challenges such as Illumination Variation, Deformation, Rotation, Aspect Ratio Change, and Partial Occlusion. Specifically, in mantis-2, while some trackers manage to track the mantis, HIPTrack-MLS consistently demonstrates superior visual performance. In person-9, due to background clutter and multiple object interference, only HIPTrack-MLS, DropTrack, ARTrack, and SeqTrack can track the target person. In phasmatodea-10, many trackers are affected by the background, leading to inaccurate tracking of the camouflaged phasmatodea. These qualitative results validate the superiority of our method for tracking camouflaged objects.
5.3 Ablation Study
Impact of the importance coefficient in Multi-Level Feature Sensing. To examine the effect of the weight coefficient $\gamma$ on balancing the contributions of the aggregated multi-level features and $H_{n}^{\rightarrow}$ in HIPTrack-MLS, we evaluated HIPTrack-MLS on the COTD dataset with a range of $\gamma$ settings. Specifically, $\gamma$ ranges from 0.1 to 1.0 with increments of 0.1. As $\gamma$ increases, the influence of the aggregated feature becomes more pronounced, while the reliance on $H_{n}^{\rightarrow}$ decreases. Table 2 details HIPTrack-MLS’s PRC and AUC on the COTD dataset for various $\gamma$ values. The results indicate that the optimal PRC and AUC, which are 52.5% and 69.2% respectively, are achieved when $\gamma$ is set to 0.1. This result indicates that an appropriate choice of $\gamma$ enhances tracking performance, confirming the effectiveness of the proposed method. Consequently, we adopted $\gamma=0.1$ as the default setting for HIPTrack-MLS.
Table 2: Illustrates the impact of the importance coefficient of aggregated feature on HIPTrack-MLS’s PRC (%) and AUC (%) on COTD. Red, blue, and green represent the first, second, and third place, respectively.
$\gamma$ 1.0 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1
PRC (%) 47.0 48.1 49.1 49.9 50.4 51.1 51.7 52.1 52.2 52.5
AUC (%) 66.7 67.4 68.0 69.1 68.1 68.2 68.7 68.8 68.3 69.2
6 Conclusion
Supplementary Material
7 Details About COTD
7.1 Annotation Rules
@@
Table 3: Co-occurrence distributions of the attributes in COTD. The diagonal demonstrates the distribution of the entire dataset, and each row or column represents the distribution of the corresponding attribute subset.
@@
Fig. 11 shows an example of applying our rules, we adhere to rule 1 by drawing bounding boxes that are tightly fitted within the boundaries of the target object, avoiding overly large sizes to ensure precision across numerous frames. Additionally, following rule 4, we focus on accurately annotating the processes of the target object’s disappearance and re-emergence (either FO or OV), which is crucial for training trackers to handle occlusions and the recapture of targets. This meticulous annotation helps create specific test scenarios, such as long-term occlusions, which are very useful for evaluating the recapture capabilities of tracking algorithms.
7.2 The Co-occurrence Distributions of the 11 Attributes in COTD
Based on the 11 attributes, the co-occurrence matrix of the COTD dataset is provided in Table 3. The horizontal and vertical axes correspond to each attribute, and the numbers in the table refer to the number of sequences with both horizontal and vertical attributes. For example, the number of sequences with both MB and DEF is 32.
(a) Tight bounding box on the first
frame.
(b) The fish is partially occluded by the hand.
(c) The fish is fully occluded by the hand.
(d) The fish partially passed the hand.
Figure 12: Benchmark results of 20 trackers and ours on COTD test set using a normalized precision plot.Best viewed by zooming in.

### v2

@@ Abstract
Index Terms:
Visual tracking; Camouflaged object tracking; Benchmark
I Introduction
@@
II Related Works
II-A Visual Tracking Algorithms
DCF-based trackers treat the visual tracking problem as an online regression task, leveraging the Parseval theorem and Fast Fourier Transform (FFT) for efficient computation. The MOSSE (Minimum Output Sum of Squared Error) filter [35] is an early representative in this domain. Since its introduction, significant advancements have been made. For example, [36, 37] introduced additional scale filters to address variations in object scale, while [38, 39] employed regularization techniques to enhance tracker robustness. Furthermore, extensions of the DCF framework, such as translation-invariant techniques, continue to evolve [40]. However, due to the limited representational capacity of handcrafted features, DCF-based trackers still face challenges in robustness under complex conditions.
To overcome the limitations of DCF-based trackers, deep learning-based trackers have garnered significant attention in recent years. These trackers utilize neural networks to automatically learn features, substantially improving tracking performance. SiamFC [41] pioneered this field by applying Siamese networks [42] to visual tracking. SiamFC learns a nonlinear similarity function through offline training and performs correlation operations on candidate regions and the initial object state to generate response maps for the search area. Building on this, the SiamRPN algorithm introduced a candidate region generation network structure with classification and regression branches to enhance tracking speed. Subsequent research, such as SiamRPN++ [43], SiamR-CNN [44], SiamCAR [45], and Ocean [46], has further improved the accuracy and robustness of Siamese trackers by incorporating techniques such as attention mechanisms, backbone network optimization, and enhanced contextual information.
Recently, the emergence of Vision Transformers (ViTs) [47] has provided new directions for the development of visual tracking technologies. Vision Transformers demonstrate potential in simplifying and unifying tracking frameworks [47, 48, 49, 50]. For instance, OStrack [50]integrates feature extraction and relational modeling, and introduces an early elimination module to achieve an efficient single-stream tracking framework. DropTrack [49] improves temporal correspondence learning by adaptively performing spatial attention dropout during frame reconstruction. HIPTrack [51] introduces a history prompting network that utilizes refined historical foreground masks and object visual features, providing comprehensive and precise guidance for trackers.
II-B Visual Tracking Benchmarks
Over the past decade, numerous benchmarks have been proposed for testing and evaluating single-object tracking performance. The advancement of visual tracking technology relies on accurate measurement and evaluation of tracking methods, with benchmark datasets serving as a critical evaluation standard. Currently, there are two main types of benchmark datasets: general benchmarks and specific benchmarks [52].
@@ II-B Visual Tracking Benchmarks
General benchmarks: General benchmark datasets are primarily used to assess the performance of tracking algorithms in typical scenarios. OTB [53, 54] and VOT [55, 56, 57, 58, 59, 60] are among the initial platforms for unified testing data and evaluation standards for general visual tracking. NFS [61] focuses on high-frame-rate videos. Many large-scale benchmarks have been introduced for training deep learning-based trackers. For example, TrackingNet [62] aims to provide resources for training and evaluating general object tracking algorithms, containing over 30,000 sequences. GOT-10K [63] includes over 10,000 videos from the WordNet semantic hierarchy to develop the generalization capability of trackers. LaSOT [64] encompasses 1,400 long-term video sequences, with an additional 150 sequences and a new evaluation scheme.
Specific benchmarks: Specific benchmark datasets aim to evaluate tracking performance for particular types of objects. For example, UAV123 [65] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [66] and PTB [67] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [52] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [68] has been introduced for tracking small and fast-moving objects, while TRO [69] focuses on tracking reflected objects.
@@
II-C Dealing with camouflaged Objects in Vision
Camouflaged objects are difficult to distinguish from their backgrounds because they blend into their surroundings. Unlike regular objects, they have similar textures, colors, and patterns as their environment, creating visual ambiguity. Recently, many computer vision techniques have been developed to tackle these challenges. For instance, Fan et al. [28] introduced the task of Camouflaged Object Detection (COD) and established a dataset COD10K with 10,000 images, along with a simple yet effective framework SINet for efficient and precise detection of camouflaged objects. Liu et al. [29] proposed a scale-adaptive correlation filter algorithm that integrates FHOG and LBP features for Camouflaged Object Segmentation (COS). Cheng et al. [22] developed a Video Camouflaged Object Detection (VCOD) framework that captures inter-frame motion and uses a spatio-temporal transformer for consistency, along with the MoCA-Mask dataset providing detailed pixel-level annotations. Hu et al. [70] tackled visual camouflage challenges in UAV surveillance with a novel trajectory planning control framework. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on consistently identifying camouflaged objects in videos while adapting to changes in position, scale, and appearance, remains under-explored. The absence of publicly available datasets for camouflaged object tracking has further constrained the progress of research and algorithm development in this field. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), specifically designed to support the development and evaluation of algorithms tailored for tracking camouflaged objects.
III Construction of the COTD Dataset
III-A Data Collection
@@
In the video selection process, we apply stringent criteria to ensure the high quality and practicality of the data. Specifically, videos are evaluated and selected based on three main criteria: 1) the diversity and complexity of the environment, 2) the frequency with which targets cross frame boundaries, and 3) the overall length and continuity of the videos. We prioritize videos that present additional challenges, such as a high degree of similarity between targets and backgrounds. Due to resource constraints in collecting, processing, and annotating video data, we select 200 sequences for the benchmark test. This ensures that each video is of high quality with comprehensive annotations, adhering to our available time, budget, and manpower constraints. The COTD dataset comprises 20 categories, namely: bird, bug, butterfly, cat, crab, dog, fish, frog, grasshopper, leopard, lizard, mantis, octopus, person, phasmatodea, seahorse, rabbit, shrimp, snake, and spider. To visually represent the composition of the dataset, we create a chart illustrating the number of sequences per category. Additionally, we generate a series of graphs depicting the maximum, minimum, and average number of frames for each category. These visual aids not only enhance the understanding of the dataset’s diversity but also highlight the challenges in tracking camouflaged objects across different scenarios. Fig. 3 presents the distribution of sequences across the 20 categories, while Fig. 4 provides a detailed view of the frame characteristics for each category, such as the range and average number of frames. These figures are crucial in demonstrating the dataset’s comprehensiveness and its potential as a robust benchmark for evaluating the performance of camouflaged object tracking algorithms. Additionally, to highlight the challenges associated with tracking camouflaged objects, we evaluate seven state-of-the-art trackers on the COTD dataset and compare them with two popular generic object tracking benchmarks. The AUC results are presented in Table I. Notably, the AUC of these trackers on the TrackingNet benchmark is more than 15.0% higher than on the COTD dataset, underscoring the significant challenges posed by tracking camouflaged objects.
@@
TABLE I: AUC (%) Comparison of state-of-the-art trackers on the COTD and generic object tracking benchmarks.
@@
TABLE II: Comparison of current popular benchmarks for visual single object tracking in the literature. $\ast$ indicates that only the test set of each dataset is reported.
Benchmark Videos Total frames Object classes Attr. Resolution Feature
OTB100 100 81K 16 11 128$\times$96 $\sim$ 800$\times$336 short-term
LaSOT* 280 685K 85 14 202$\times$360 $\sim$ 1280$\times$720 category balance
GOT-10k* 420 56K 84 6 270$\times$480 $\sim$ 3840$\times$2160 generic
VOT 62 20K 37 9 320$\times$180 $\sim$ 1280$\times$720 annual
COTD 200 80K 29 12 244$\times$578 $\sim$ 2400$\times$1080 camouflaged
By contrast, our proposed COTD is the first benchmark dataset to focus on tracking camouflaged objects. A summarized comparison with existing benchmarks is reported in Table II.
(a) Tight bounding box on the first
frame.
(b) The fish is partially occluded by the hand.
(c) The fish is fully occluded by the hand.
(d) The fish partially passed the hand.
III-B Data Annotation
We follow the principles in [62] for sequence annotation. Each frame undergoes meticulous annotation by expert annotators, comprising students specializing in visual tracking research. The annotation process involves drawing axis-aligned bounding boxes that represent the most compact boundary of the target object, with careful documentation of object visibility, occlusion, and absence states. Examples of bounding box annotations in the COTD are shown in Fig. 2.
To ensure annotation quality and consistency, we implement a rigorous two-stage verification process. In the primary stage, annotators conduct frame-by-frame labeling, ensuring precise object identification even in complex scenes. This is followed by a secondary verification where another annotator reviews all annotations, with any discrepancies resolved through discussion and consensus.
@@ III-B Data Annotation
As illustrated in Fig. 5, we adhere to the rules by drawing bounding boxes that are tightly fitted within the boundaries of the target object, avoiding overly large sizes to ensure precision across numerous frames. Additionally, we focus on accurately annotating the processes of the target object’s disappearance and re-emergence (either FO or OV), which is crucial for training trackers to handle occlusions and the recapture of targets. This meticulous annotation helps create specific test scenarios, such as long-term occlusions, which are very useful for evaluating the recapture capabilities of tracking algorithms.
III-C Attributes
Figure 6: Distribution of sequences with each attribute in COTD.
Figure 8: Structure of the Multi-Level Feature Sensing Encoder.
Following the prevalent tracking benchmark methodologies [78], [63], and [54], we meticulously annotate each video sequence with detailed attributes to enable an in-depth analysis of tracking algorithms’ performance. The proposed COTD dataset features 12 attributes, with each sequence including Background Clutter (BC) to highlight the dataset’s camouflage characteristics. In addition, we annotate the following 11 attributes: (1) Illumination Variation (IV), (2) Scale Variation (SV), (3) Deformation (DEF), (4) Motion Blur (MB), (5) Fast Motion (FM), which is marked when the object center moves by at least 20% of its own size between consecutive frames, (6) Out-of-View (OV), (7) Low Resolution (LR), marked when the object area is smaller than 900 pixels, (8) Partial Occlusion (POC), (9) Rotation (ROT), (10) Full Occlusion (FOC), and (11) Aspect Ratio Change (ARC), marked when the object’s bounding box aspect ratio falls outside the range [0.5, 2]. Fig.6 shows the distribution of these attributes on COTD. As can be seen, the most common challenge in COTD is Motion Blur. In addition, the Deformation and Partial Occlusion also present frequently in COTD.
TABLE III: Co-occurrence distributions of the attributes in COTD. The diagonal demonstrates the distribution of the entire dataset, and each row or column represents the distribution of the corresponding attribute subset.
@@ III-C Attributes
Based on the 11 attributes, the co-occurrence matrix of the COTD dataset is provided in Table III. The horizontal and vertical axes correspond to each attribute, and the numbers in the table refer to the number of sequences with both horizontal and vertical attributes. For example, the number of sequences with both MB and DEF is 32. By comprehensively considering these attributes, we can conduct a thorough evaluation of tracking algorithms under various challenging conditions.
IV Method
IV-A Multi-Level Feature Sensing Encoder
IV-B Overall Architecture
V Evaluation
In this study, precision and success rate serve as the quantitative measures for evaluation.
Precision plot: The precision plot is a standard metric used for evaluating tracking performance. It assesses tracking accuracy by measuring the center position error, which calculates the average Euclidean distance between the tracked object’s center and its ground truth position. This metric summarizes the tracking performance by averaging the errors across all frames in a sequence. However, when a tracker loses the object, its output may become arbitrary, reducing the reliability of the average error [80]. To mitigate this issue, the precision plot is commonly used. It evaluates the overall tracking accuracy by comparing the tracker’s estimated position to the ground truth within a specified distance threshold [80, 81]. In our evaluation, we use a 20-pixel threshold for the precision score, in line with established practices [80].
Success plot: Another evaluation metric commonly used is the success plot. This metric gauges tracker performance by analyzing the overlap between predicted and ground truth bounding boxes. Specifically, for a given tracking bounding box $r_{t}$ and a ground truth bounding box $r_{a}$ , the overlap score is computed as $S=\frac{\left|r_{t}\bigcap r_{a}\right|}{\left|r_{t}\bigcup r_{a}\right|}$ , where $\cap$ and $\cup$ denote the intersection and union of the two bounding boxes, respectively, and ||$\cdot$|| ⋅ | represents the number of pixels contained within these regions. To assess tracking performance across a sequence of frames, we count the number of frames where the overlap score $S$ exceeds a predefined threshold $t_{o}$. The success plot then illustrates the proportion of frames deemed successful at various overlap thresholds ranging from 0 to 1. This visual representation helps in understanding how often the tracker achieves sufficient overlap with the ground truth over different levels of thresholding.Relying on a single success rate at a specific threshold (e.g. $t_{o}$=0.5) may not offer a comprehensive view of a tracker’s performance, as it provides a limited perspective on tracking accuracy. To address this, the area under the curve (AUC) is typically employed in conjunction with the success plot to rank tracking algorithms [80, 81]. The AUC aggregates the success rate across all threshold values, offering a thorough evaluation of the tracking algorithm’s performance across various conditions. This method facilitates a more nuanced and accurate comparison of different tracking algorithms by providing a holistic view of their effectiveness in various scenarios. Each tracker is evaluated using the default parameters provided in the source code. We employ one-pass evaluation (OPE) and assess precision and success rate, as per [82, 83].
V-A Trackers for Comparison
V-B Evaluation Results
Figure 11: Qualitative evaluation on 4 sequences from COTD, i.e. fish-4, mantis-2, person-9, and phasmatodea-10 from top to bottom. The results of different methods have been shown with different colors, and ‘GT’ denotes the ground truth.
Attribute-based performance: We conducte a comprehensive performance evaluation on eleven common attributes to better analyze and understand the performance of different trackers on the COTD dataset. HIPTrack-MLS consistently ranks first across most attribute subsets. Due to paper length constraints, we present the PRC results for four common challenges in Fig. 10, including Aspect Ration Change (ARC), Illumination Variation (IV), Out-of-View (OV), and Partial Occlusion (POC). As shown, HIPTrack-MLS consistently ranks first across all these attribute subsets, surpassing the baseline HIPTrack by at least 1.0% in all these subsets. Notably, HIPTrack-MLS achieves a precision of 0.271 in the IV subset, surpassing the second place HIPTrack by 3.4%. For the ARC and OV attributes, HIPTrack-MLS shows improvements of 1.3% and 1.0%, respectively, over HIPTrack.